The height (t) in feet above the ground, of a ball thrown into the air from the top of a building is given by the function of h(t)= -16(t-2)^2+264, where t is the number of seconds the ball has been in the air.
a. From what height was the ball thrown?
b. what height does the ball reach?
c. when does the ball strike the ground?
e. for what values of t does the function have meaning?

Answer by Alan3354(69443)

The height (t) in feet above the ground, of a ball thrown into the air from the top of a building is given by the function of h(t)= -16(t-2)^2+264, where t is the number of seconds the ball has been in the air.
a. From what height was the ball thrown?
b. what height does the ball reach?
c. when does the ball strike the ground?
e. for what values of t does the function have meaning?
------------
I think h is the height, not t.
h(t)= -16(t-2)^2+264
h(t) = -16t^2 + 64t + 200
---------------------
a. From what height was the ball thrown?
200 feet
b. what height does the ball reach?
264 feet
c. when does the ball strike the ground?
When h = 0
-16t^2 + 64t + 200 = 0
t = 6.062 seconds
e. for what values of t does the function have meaning?
From launch to impact, t = 0 to t = 6.062 seconds
Later than impact, the ball is stationary, h = 0. The meaning is a matter of opinion.